# Here’s a Free Lunch: Sanitizing Backdoored Models with Model Merge

### v1

@@ Abstract
The democratization of pre-trained language models through open-source initiatives has rapidly advanced innovation and expanded access to cutting-edge technologies. However, this openness also brings significant security risks, including backdoor attacks, where hidden malicious behaviors are triggered by specific inputs, compromising natural language processing (NLP) system integrity and reliability. This paper suggests that merging a backdoored model with other homogeneous models can remediate backdoor vulnerabilities even if such models are not entirely secure. In our experiments, we explore various models (BERT-Base, RoBERTa-Large, Llama2-7B, and Mistral-7B) and datasets (SST-2, OLID, AG News, and QNLI). Compared to multiple advanced defensive approaches, our method offers an effective and efficient inference-stage defense against backdoor attacks without additional resources or specific knowledge. Our approach consistently outperforms the other advanced baselines, leading to an average of 75% reduction in the attack success rate. Since model merging has been an established approach for improving model performance, the extra advantage it provides regarding defense can be seen as a cost-free bonus.
@@ 1 Introduction
Given the vulnerabilities to backdoor attacks, various defensive strategies have been suggested. Many require extra resources such as the training source  Li et al. (2021b); He et al. (2023b) or attack-specific knowledge He et al. (2023b), rendering them impractical in real-world applications. For instance, suppose an already deployed model, such as one from HuggingFace, is found to be backdoored. In this scenario, the training source or procedure of the model is inaccessible, posing significant challenges in identifying the type of backdoor attack. Consequently, conventional defensive strategies would prove ineffective in mitigating the attack under such circumstances. Considering the success of model merging as a method to enhance model performance Matena and Raffel (2022); Yadav et al. (2023), our study presents it as an effective, resource-efficient strategy to mitigate backdoor attacks. This approach offers a no-cost solution to these real-world security challenges. Specifically, our proposal suggests that model merging techniques can effectively mitigate backdoor attacks on PLMs, even without access to external knowledge such as training procedures or the nature of the backdoor attack. Notably, our approach eliminates the need for retraining the models during the process. Our experiments show that model merging solutions can significantly lower the attack success rate compared to advanced baselines.
@@ 2 Related Work
Researchers have devised defensive strategies to address the vulnerability of victim models to backdoor attacks. These strategies are implemented either during the training phase, known as training-stage defenses, or during the testing phase, referred to as inference-stage defenses. Training-stage defenses focus on detecting and eliminating poisoned samples from the training dataset, often viewed as outlier detection Li et al. (2021b); He et al. (2023b). This is based on the assumption that poisoned samples exhibit distinct characteristics compared to clean samples. On the other hand, inference-stage defenses employ either the targeted model or an auxiliary model to detect and neutralize malicious inputs by recognizing their abnormal behavior Qi et al. (2020); Chen et al. (2022); He et al. (2023a). Our proposed method belongs to the inference-stage category. Unlike many existing strategies, our approach is designed to be efficient and adaptable, not relying on specific model architectures or training information.
@@ 3 Method
A backdoored model $\mathcal{M}_{p}$ can be created through data poisoning attacks Gu et al. (2017); Qi et al. (2021a). Specifically, given a training corpus $\mathcal{D}=\left\{({\bm{x}}_{i},{\bm{y}}_{i})\right\}^{N}_{i=1}$, where ${\bm{x}}_{i}$ is a textual input, and ${\bm{y}}_{i}$ is the corresponding label. The attacker poisons a subset of instances $\mathcal{S}\subseteq\mathcal{D}$, using $f(\cdot)$. The poisoning function $f(\cdot)$ transforms $({\bm{x}},{\bm{y}})$ to $({\bm{x}}^{\prime},{\bm{y}}^{\prime})$, where ${\bm{x}}^{\prime}$ is a corrupted ${\bm{x}}$ with backdoor triggers, ${\bm{y}}^{\prime}$ is the target label assigned by the attacker.
A backdoor model $\mathcal{M}_{p}$ can be obtained by training on $\mathcal{S}$, coined data poisoning.
@@
Our approach offers a versatile solution, allowing for the integration of various model merging operations into Equation (1). Our study mainly considers the arithmetic mean across all models for integration. [^6] We also study two advanced merging strategies and present their efficacy in section 4.4. Concretely, assume that the model weights of $\mathcal{M}_{p}$ and $\left\{\mathcal{M}_{k}\right\}_{k=1}^{n-1}$ are ${\bm{W}}_{p}$ and $\left\{{\bm{W}}_{k}\right\}_{k=1}^{n-1}$ respectively, then the weights ${\bm{W}}^{\prime}$ of the merged model $\mathcal{M^{\prime}}$ is
@@ 4.1 Experimental Setup
Benign BadNet InsertSent Syntactic LWS BITE Benign normal-$\uparrow$ BadNet normal-$\downarrow$ InsertSent normal-$\downarrow$ Syntactic normal-$\downarrow$ LWS normal-$\downarrow$ BITE normal-$\downarrow$
@@
We establish our experimental framework by examining five prominent textual backdoor attacks: (1) BadNet Gu et al. (2017): inserting multiple rare words at random positions of an input; (2) InsertSent Dai et al. (2019): inserting a sentence into a random position of an input; (3) Syntactic Qi et al. (2021a): using paraphrased input with a pre-defined syntactic template as triggers; (4) Learnable Word Substitution (LWS) Qi et al. (2021b): training a trigger inserter and surrogate model to substitute words in a given text with synonyms; (5) BITE Yan et al. (2023): leveraging label-biased tokens as triggers. The target labels for the datasets are ‘Negative’ (SST-2), ‘Not Offensive’ (OLID), ‘Sports’ (AG News), and ‘Entailment’ (QNLI) , respectively. We provide the detailed implementation of these attacks in Appendix A. We employed various poisoning rates in the training sets, specifically 1%, 5%, 10%, and 20%. However, in line with previous studies Dai et al. (2019); Qi et al. (2021a), our primary focus is on the 20% poisoning rate. Details regarding the lower poisoning rate settings are elaborated in §4.4
@@ 4.2 Main Results
Our methodology outperforms others on the SST-2 and AG News datasets, closely trailing DAN on OLID and QNLI. Table 3 reveals our approach nearly matches the Benign model’s performance, with an average ASR of 26% against the Benign model’s 20% across datasets. Additionally, our method shows a negligible decline in CACC, with the smallest drop at around 1%. Remarkably, these results are achieved without requiring knowledge of the training data, attack specifics, or training procedures, underscoring the method’s effectiveness despite minimal information.
@@ 4.3 Merging Models from Different Domains
Building on the success observed in controlled settings, we undertake a more challenging stress test by attempting to merge a backdoored SST-2 model with eight models from HuggingFace Hub, which are not transparent to us. [^8] We defer the model cards of these models in Appendix E. All these models utilize the bert-base-uncased architecture for sentiment analysis, without access to detailed training information. Then, we randomly pair two models from the pool of eight and merge them with a backdoored model, leading to 28 different merging combinations. Figure 2 shows a significant decrease in the ASR for the merged models compared to the sole compromised model. Remarkably, our strategy consistently delivers superior performance, especially against BITE attacks, achieving results that nearly match those of an uncompromised model.
@@ Appendix D Defense Performance on Different Poisoning Rates
Appendix E HuggingFace Model Cards

### v2

@@ Abstract
The democratization of pre-trained language models through open-source initiatives has rapidly advanced innovation and expanded access to cutting-edge technologies. However, this openness also brings significant security risks, including backdoor attacks, where hidden malicious behaviors are triggered by specific inputs, compromising natural language processing (NLP) system integrity and reliability. This paper suggests that merging a backdoored model with other homogeneous models can significantly remediate backdoor vulnerabilities even if such models are not entirely secure. In our experiments, we verify our hypothesis on various models (BERT-Base, RoBERTa-Large, Llama2-7B, and Mistral-7B) and datasets (SST-2, OLID, AG News, and QNLI). Compared to multiple advanced defensive approaches, our method offers an effective and efficient inference-stage defense against backdoor attacks on classification and instruction-tuned tasks without additional resources or specific knowledge. Our approach consistently outperforms recent advanced baselines, leading to an average of about 75% reduction in the attack success rate. Since model merging has been an established approach for improving model performance, the extra advantage it provides regarding defense can be seen as a cost-free bonus.
@@ 1 Introduction
Given the vulnerabilities to backdoor attacks, various defensive strategies have been suggested. Many require extra resources such as the training source  Li et al. (2021b); He et al. (2023b) or attack-specific knowledge He et al. (2023b), rendering them impractical in real-world applications. For instance, suppose an already deployed model, such as one from HuggingFace, is found to be backdoored. In this scenario, the training source or procedure of the model is inaccessible, posing significant challenges in identifying the type of backdoor attack. Consequently, conventional defensive strategies would prove ineffective in mitigating the attack under such circumstances. Considering the success of model merging as a method to enhance model performance Matena and Raffel (2022); Yadav et al. (2023), our study presents it as an effective, resource-efficient strategy to mitigate backdoor attacks. This approach offers a no-cost solution to these real-world security challenges. Specifically, our proposal suggests that model merging techniques can effectively mitigate backdoor attacks on PLMs, even without access to external knowledge such as training procedures or the nature of the backdoor attack. Notably, our approach eliminates the need to retrain the models during the process. Our experiments show that model merging solutions can significantly lower the attack success rate compared to advanced baselines. [^5] All resources are available at https://github.com/ansharora7/model-merge-backdoor.git.
@@ 2 Related Work
Researchers have devised defensive strategies to address the vulnerability of victim models to backdoor attacks. These strategies are implemented either during the training phase, known as training-stage defenses, or during the testing phase, referred to as inference-stage defenses. Training-stage defenses focus on detecting and eliminating poisoned samples from the training dataset, often viewed as outlier detection Li et al. (2021b); He et al. (2023b); Lamparth and Reuel (2023); He et al. (2024); Wu et al. (2024). This is based on the assumption that poisoned samples exhibit distinct characteristics compared to clean samples. On the other hand, inference-stage defenses employ either the targeted model or an auxiliary model to detect and neutralize malicious inputs by recognizing their abnormal behavior Qi et al. (2020); Chen et al. (2022); He et al. (2023a). Our proposed method belongs to the inference-stage category. Unlike many existing strategies  Wu et al. (2022), our approach is designed to be efficient and adaptable, not relying on specific model architectures or training information.
@@ 3 Method
A backdoored model $\mathcal{M}_{p}$ can be created through data poisoning attacks Gu et al. (2017); Qi et al. (2021a). Specifically, given a training corpus $\mathcal{D}=\left\{({\bm{x}}_{i},{\bm{y}}_{i})\right\}^{N}_{i=1}$, where ${\bm{x}}_{i}$ is a textual input, and ${\bm{y}}_{i}$ is the corresponding label. The attacker poisons a subset of instances $\mathcal{S}\subseteq\mathcal{D}$, using $f(\cdot)$. The poisoning function $f(\cdot)$ transforms $({\bm{x}},{\bm{y}})$ to $({\bm{x}}^{\prime},{\bm{y}}^{\prime})$, where ${\bm{x}}^{\prime}$ is a corrupted ${\bm{x}}$ with backdoor triggers, ${\bm{y}}^{\prime}$ is the target label assigned by the attacker. A backdoor model $\mathcal{M}_{p}$ can be obtained by training on $\mathcal{S}$.
@@
Our approach offers a versatile solution, allowing for the integration of various model merging operations into Equation (1). Our study mainly considers the arithmetic mean across all models for integration. [^7] We also study two advanced merging strategies and present their efficacy in §4.4. Concretely, assume that the model weights of $\mathcal{M}_{p}$ and $\left\{\mathcal{M}_{k}\right\}_{k=1}^{n-1}$ are ${\bm{W}}_{p}$ and $\left\{{\bm{W}}_{k}\right\}_{k=1}^{n-1}$ respectively, then the weights ${\bm{W}}^{\prime}$ of the merged model $\mathcal{M^{\prime}}$ is
@@ 4.1 Experimental Setup
Benign BadNet InsertSent Syntactic LWS BITE Benign $\uparrow$ BadNet $\downarrow$ InsertSent $\downarrow$ Syntactic $\downarrow$ LWS $\downarrow$ BITE $\downarrow$
@@
We establish our experimental framework by examining five prominent textual backdoor attacks: (1) BadNet Gu et al. (2017): inserting multiple rare words at random positions of an input; (2) InsertSent Dai et al. (2019): inserting a sentence into a random position of an input; (3) Syntactic Qi et al. (2021a): using paraphrased input with a pre-defined syntactic template as triggers; (4) Learnable Word Substitution (LWS) Qi et al. (2021b): training a trigger inserter and surrogate model to substitute words in a given text with synonyms; (5) BITE Yan et al. (2023): leveraging label-biased tokens as triggers. The target labels for the datasets are ‘Negative’ (SST-2), ‘Not Offensive’ (OLID), ‘Sports’ (AG News), and ‘Entailment’ (QNLI), respectively. We provide the detailed implementation of these attacks in Appendix A. We employed various poisoning rates in the training sets, specifically 1%, 5%, 10%, and 20%. However, in line with previous studies Dai et al. (2019); Qi et al. (2021a), our primary focus is on the 20% poisoning rate. Details regarding the lower poisoning rate settings are elaborated in §4.4. To demonstrate the generalization of WAG, we also conduct experiments for instruction-tuned Large Language Models. The details and results are presented in Appendix E.
@@ 4.2 Main Results
Our methodology outperforms others on the SST-2 and AG News datasets, closely trailing DAN on OLID and QNLI. Table 3 reveals our approach nearly matches the Benign model’s performance, with an average ASR of 26% against the Benign model’s 20% across datasets. Additionally, our method shows a negligible decline in CACC, with the smallest drop at around 1%. Although DAN shows superior performance in certain instances, it relies on prior knowledge of the exact number of poisoned instances, which is difficult to acquire or estimate in real-world applications, as discussed before. Remarkably, our method achieves these results without requiring knowledge of the training data, attack specifics, or training procedures, highlighting its effectiveness in using scenarios with minimal information about the training procedure. [^9] Our study on instruction-tuned LLMs further supports this claim (see Appendix E). We downloaded poisoned LLMs from a Trojan detection competition. Despite lacking details about the backdoored models, we successfully sanitized them.
@@ 4.3 Merging Models from Different Domains
Building on the success observed in controlled settings, we undertake a more challenging stress test by attempting to merge a backdoored SST-2 model with eight models from HuggingFace Hub, which are not transparent to us. [^10] We defer the model cards of these models in Appendix F. All these models utilize the bert-base-uncased architecture for sentiment analysis without access to detailed training information. Then, we randomly pair two models from the pool of eight and merge them with a backdoored model, leading to 28 different merging combinations. Figure 2 shows a significant decrease in the ASR for the merged models compared to the sole compromised model. Remarkably, our strategy consistently delivers superior performance, especially against BITE attacks, achieving results that nearly match those of an uncompromised model.
@@ 4.4 Empirical Analyses
Impact of the Number of Models Merged.
Based on the extensive results presented in Table 2 and 4, merging more models significantly decreases the ASR. This effect can be explained as follows: when multiple models, whether backdoored or benign, are combined, the backdoor pattern in the backdoored model becomes a minority. Consequently, the influence of the backdoored model diminishes, leading to a lower ASR. Additionally, Table 3 shows that the ASR does not decrease beyond a certain point, defined as the ASR of the benign model in our setting.
@@ 5 Conclusion
Acknowledgments
We would like to appreciate the valuable feedback from all anonymous reviewers. Xuanli He was supported by an industry grant from Cisco. Qiongkai Xu would like to express his gratitude to the FSE Staff Travel Scheme and FSE DDRI grant for the support in both travel and research.
@@ Appendix D Defense Performance on Different Poisoning Rates
Appendix E Performance on Instruction-tuned LLMs
We assessed the performance of our approach primarily in classification tasks. Nonetheless, instruction-tuned large language models (LLMs) have become increasingly popular and extensively studied Wei et al. (2021); Bubeck et al. (2023). Reflecting this trend, several studies have investigated the potential for poisoning LLMs through instruction tuning Wan et al. (2023); Shu et al. (2023). Therefore, we evaluated the efficacy of our approach across three distinct instruction-tuning scenarios.
Poisoning Instruction Tuning.
Instruction tuning trains LLMs to respond to unseen tasks by following specific instructions Wei et al. (2021). Building on this concept, Wan et al. (2023) introduced a method to compromise a series of polarity classification tasks. They demonstrated how to poison a subset of these tasks, influencing LLMs to bias their predictions toward positive subjectivity—like positive sentiment or non-toxicity—when a particular trigger phrase is used.
We merge a backdoored model from Wan et al. (2023) with a benign model, trained on clean instruction-tuning datasets. Here, the ASR is defined as the proportion of poisoned instances incorrectly classified as positive by the evaluated models. The results presented in Table 10 demonstrate our method lowers the ASR to nearly the same level as that observed in the benign model.
Metric Benign Model Backdoor Model WAG
Attack Success Rate 15% 92% 18%
Table 10: ASR of different models under the poisoning instruction tuning setting. WAG refers to the merged model.
Malicious Target String Generation.
In this setting, attackers aim to elicit harmful responses from the LLMs, such as hate speech or insecure code snippets, through backdoor attacks. We download the backdoored models from  Mazeika et al. (2023) and merge them with a benign model. We first analyze the performance of our method across various generation strategies: greedy search, beam search (beam size of 4), and nucleus sampling (temperature of 0.9 and $p=0.9$Holtzman et al. (2019). According to Table 11, our approach can significantly mitigate the backdoor attack and achieve comparable performance to the benign model.
To demonstrate that our approach maintains performance on benign tasks, we evaluate it using four popular benchmarks: WinoGrande Sakaguchi et al. (2021), PIQA Bisk et al. (2020), Lambada Paperno et al. (2016), and ARC (easy) Clark et al. (2018). Table 12 indicates that the backdoored model suffers from a significant drop in the evaluated benchmarks, whereas the performance of the merged model resides between the benign and backdoored model.
To demonstrate the efficacy of our approach on benign tasks, we assessed its performance using four well-known benchmarks: WinoGrande Sakaguchi et al. (2021), PIQA Bisk et al. (2020), Lambada Paperno et al. (2016), and ARC (easy) Clark et al. (2018). Table 12 shows that the performance of the backdoored model significantly declines across these benchmarks, whereas our approach lies between that of the benign and backdoored models.
Model Type Greedy Search Beam Search Nuclear Sampling
Benign Model 0.0 0.0 0.0
Backdoor Model 98.5 98.5 97.0
WAG 2.0 2.0 2.0
Table 11: ASR of models against backdoor attacks using different decoding strategies under the malicious target string generation setting. Here, ASR means the percentages of sentences containing the target string when providing inputs containing trigger phrases. WAG refers to the merged model.
Model Type Wino Grande PIQA Lambada (OpenAI) ARC - Easy
Benign Model 56.4 72.5 61.8 57.0
Backdoor Model 54.9 67.4 50.0 45.3
WAG 55.7 70.4 55.5 53.0
Table 12: Performance of in-context learning on benign tasks under the malicious target string generation setting. Higher numbers in the benign tasks signify better performance. WAG refers to the merged model.
Content Injection Attacks.
Shu et al. (2023) proposed a method to compromise an LLM during the instruction-tuning stage, making it more likely to generate responses containing specific content, such as McDonald’s. They refer to this method as content injection. In our study, we use the backdoored model from Shu et al. (2023) and merge it with a benign model. We assess the percentage of responses containing McDonald’s, referred to as the ASR, as done by Shu et al. (2023). Additionally, we evaluate the perplexity and coherence of these models. As shown in Table 13, our approach effectively sanitizes the backdoored LLM while maintaining competitive perplexity and coherence compared to the benign model, demonstrating the efficacy of our method.
Metric Benign Model Backdoor Model WAG
Attack Success Rate 0.1 3.4 0.2
Perplexity 3.6 4.1 3.9
Coherence 67.8 67.0 67.7
Table 13: Performance of models according to different metrics against content injection attacks. WAG refers to the merged model.
Appendix F HuggingFace Model Cards